An investment of $4600 is made at an annual simple interest rate of 6.8%. How much additional money must be invested at an annual simple interest rate of 9% so that the total interest earned is 8% of the total investment?
Answer by checkley77(12844)

4,600*.068+.09X=.08(4,600+X)
312.8+.09X=368+.08X
.09X-.08X=368-312.8
.01X=55.2
X=55.2/.01
X=5,520 IS THE ADDITIONAL INVESTMENT REQUIRED.
PROOF:
4,600*.068+.09*5,520=.08(4,600+5,520)
312.8+496.8=.08*10,120
809.6=809.6